OpenStudy (anonymous):

a= 2 (b+c) solve for c

OpenStudy (anonymous):

Ready to work on this together? Just lemme know when you're ready. :P

OpenStudy (anonymous):

yes I am

OpenStudy (anonymous):

Awesome. Let's do it then :D

OpenStudy (anonymous):

First problem is that our C is stuck inside a parenthesis. How do we help it escape the parenthesis?

OpenStudy (anonymous):

(Don't worry about your other questions yet. We'll work on those when we finish this one)

OpenStudy (anonymous):

ok use distribute?

OpenStudy (anonymous):

Yup! So what do you get?

OpenStudy (anonymous):

a= 2b + 2c

OpenStudy (anonymous):

a= 2 (b+c) a = 2b + 2c Good! So now what do we do to further get the C closer to being by itself?

OpenStudy (anonymous):

Hint: There's a 2b and a 2 attached to the C. Is it easier to divide stuff? Or to subtract stuff?

OpenStudy (anonymous):

divide

OpenStudy (anonymous):

To be honest, I think subtraction's easier. BUT, it works both ways. So what are we left with after dividing by 2?

OpenStudy (anonymous):

Whoops... gave away the number. But you were going to divide by 2 anyway, right?

OpenStudy (anonymous):

so it cancels out right?

OpenStudy (anonymous):

Well, the 2s cancel out. But don't forget: if you divide one side by two, you must do the same to the other. But, would you like to see an easier way?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Okie dokie. When you were at this step: a = 2b + 2c It'd be easier to make the 2b go to the other side. How can we move it to there?

OpenStudy (anonymous):

Hint: if you try dividing, you'd be left with a/2b = 2c/2b+1 So... you can tell you don't want to divide.

OpenStudy (anonymous):

Still here roseanne? I can tell you what to do if it's confusing.

OpenStudy (anonymous):

no I understand

OpenStudy (anonymous):

How am I going to know that on my midterm

OpenStudy (anonymous):

im just wanting to understand

OpenStudy (anonymous):

Okie dokie. So what do we do with a = 2b + 2c to isolate the 2c further?

OpenStudy (anonymous):

subtract

OpenStudy (anonymous):

That's right :D Now that we have -2b + a = 2c, what is our final step?

OpenStudy (anonymous):

Hint: It's easier if you view it with parentheses: (-2b + a) = 2c

OpenStudy (anonymous):

So let's do that last step and go onto the other questions.

OpenStudy (anonymous):

Still here roseanne? Just tell me what the final step is and we can get crackin' on the next few questions. These will be a lot easier since you'll know the steps.